Compute the expected return given these three economic states, their likelihoods, and the potential returns: (Round your answer to 2 decimal places.)

Economic State | Probability | Return

Fast growth: 0.29 30 %

Slow growth: 0.4     3

Recession: 0.30    –27
Answer #1

The formula for expected return is:

Expected return = p1 * r1 + p2 * r2 + p3 * r3

where, p1,p2 and p3 are the probabilities and r1,r2 and r3 are the returns for economic states.

Putting the given values of the probability in the above formula, we get,

Expected return = (0.29 * 30%) + (0.4 * 30%) + (0.3 * -27%)

Expected return = 8.7 + 12 - 8.1

Expected return = 12.6%